# Network Structure Trumps Neuron Dynamics: Insights from Drosophila Connectome Simulations

### v3

@@ Abstract
Despite the success of artificial neural networks, the necessity of real network structures in simulating intelligence remains unclear. Utilizing the largest adult Drosophila connectome data set, we constructed a large-scale network communication model framework based on simple neuronal activation mechanisms to simulate the activation behavior observed in the connectome. The results demonstrate that even with simple propagation rules, models based on real neural network structures can generate activation patterns similar to those in the actual brain. Importantly, we found that using different neuronal dynamics models, all produced similar activation patterns. This consistency across different models emphasizes the crucial role of network topology in neural information processing, challenging views that rely solely on neuron count or complex individual neuron dynamics. Moreover, we test the influence of network reconnect rate and find that at least 1%’s reconnect rate will ruin the activation patterns appeared before. By comparing network distances and spatial distances, we found that network distance better explains the information propagation patterns between neurons, highlighting the importance of topological structure in neural information processing. To facilitate these studies, we developed real-time 3D large spatial network visualization software, bridging a crucial gap in existing tools. Our findings underscore the importance of network structure in neural activation and provide new insights into the fundamental principles governing brain functionality.
@@ 1 Introduction
The neural systems of intelligent beings maintain constant communication[1, 2]. The intricate mechanism of this communication unfolds as follows: specific neurons, often input neurons, receive stimulus signals from the external environment or other parts of the body, changing from a resting to an activated state. Subsequently, they relay these stimulus signals to the next tier of neurons through synapses. This process gives rise to macroscopic emergence across the entire neural system. Such communication fosters the emergence of intelligence in all intelligent life forms, including human beings. Consequently, comprehending and replicating this process has become an indispensable avenue for unraveling biological intelligence and realizing artificial intelligence[3].
Neural systems are interconnected networks[4], with neurons engaging in constant communication via synapses within the network’s topology. Some scientists think the brain’s unique and complex network structure is pivotal to the emergence of intelligence[5]. Many scientists posit that network science is crucial for deciphering the mechanisms underlying the flexible regulation of neural communication[6, 7, 8]. Determining how to leverage network science methodologies stands as a paramount challenge in contemporary neuroscience.
However, whether this network structure determines the emergence of intelligence remains a question. In recent years, artificial neural networks (ANNs) have taken rapid development in enormous fields[9, 10, 11]. A notable recent achievement is large language model (LLM). However, the structure of any ANNs does not follow the actual brain network structure strictly, which poses a challenge to the viewpoint that network structure inherently contains intelligence. Simultaneously, many scientists in machine learning adhere to the scaling law: the more parameters the network has, the greater the degree of emergent intelligence[12]. This viewpoint suggests that intelligence is solely related to the number of parameters a model has, rather than the unique network structure between neurons. Under this perspective, the reason biological brains generate intelligence is due to the number of neurons, and the unique network structure is a coincidence of evolution.
To demonstrate the importance of brain network structure, network scientists are constantly exploring and experimenting. An approach is brain network communication model[13]. Scientists wants to use brain network communication model to understand and describe the interaction and communication between different parts of the brain. This model aids in studying the structural and functional aspects of neuronal networks and how they coordinate to accomplish complex cognitive processes[14].
The first challenge of this approach is obtaining a complete and accurate mapping of the brain’s structure. For example, current imaging technologies, such as MRI, fMRI, and electron microscopy, have limitations in resolution and speed. While MRI and fMRI provide non-invasive imaging, their spatial resolution is not sufficient to map individual neurons and synapses. On the other hand, techniques like electron microscopy, which can achieve higher resolution, are extremely slow and generate enormous amounts of data that are challenging to process[15]. Modeling from the micro scale is necessary because modeling at the level of brain regions or nerve fibers does not include the most basic structural information and lacks the most complete physical foundation.
Moreover, current models often rely on prevalent metrics such as signaling cost and activation time to assess model quality[16]. Yet, these metrics fall short of verifying whether a network model accurately reflects the brain’s actual activation patterns, which is the fundamental goal of brain simulation. There is a need for evaluation criteria that can affirm whether network communication models genuinely mirror the brain’s communication processes.
Based on facts and challenges above, benefiting from a largest-to-date adult Drosophila connectome data[17], this study attempts to investigate how the type of structure of neurons could affect the activation patterns of a brain. Our approach involves crafting a large scale network communication model framework based on simple mechanisms, designed to simulate the activation behavior observed in the connectome. Specifically, we establish the network representation of the connectome. Then we design a large scale threshold propagation model of nerve cells: when a neuron received enough stimuli, it activates and subsequently transmits signals to other neurons. After a few iterations, the whole network emerges an activation pattern. We use threshold model and sigmoid model as the basic neuron model. We also design an improved LIF model as the basic neuron model for comparison. Additionally, we propose an evaluation criteria based on the ratio of actually activated neurons to the expected number of activation. Experiments reveals that, from threshold , sigmoid, and LIF neuron model, no matter activation mechanism is, our model has significant activation in the type of neurons that should be affected by the stimulus and slight activation in irrelevant neurons. Moreover, to supply more evidence of the vitalness of network structure, we disturb the network by the way of degree-preserving rewiring and and find the disturbed network can’t generate any correct activation pattern.
In addition, we test the sensitivity of disturb degree. For details, we record the activation rate of areas under different reconnect rates and find that 1%’s random change of network structure will cause a complete change in function.
To prove the network structure, rather than the spatial structure, matters in activation pattern generation, we calculate the average network and spatial distance of input and activation neurons and results reveal that the average network distance between input neurons and their corresponding activation areas is shorter than unrelated areas, which is a strong evidence of network matters in message propagation in brain.
Furthermore, when providing unilateral stimuli to input neurons, we observe a bilateral response in the Drosophila connectome, aligning with reality. Then we calculate the statistics properties and find that the left and right hemispheres have extremely similar degree, clustering coefficient, and eigenvector centrality vector coefficient. The vectors’ Pearson coefficient are up to 0.9986, 0.9988, and 0.9991, respectively.
@@
We also developed a 3D network visualization software see in appendix F, utilizing HTML5’s Canvas and Three.js based on WebGL technology for large spatial network visualization effects. This tool enables real-time observation of each neuron’s communication process.
@@ 2 Results
First, we report the basic statistics of Drosophila connectome. After that, to test the performance of our model in terms of realism on the data set and determine whether different artificial mechanisms can achieve similar activation patterns when the network structure changes, we give different stimuli and record the activation pattern of different neurons. In addition, we give a unilateral stimulus (stimulus one hemisphere) and record the bilateral response of neurons. Besides, to add more neurons’ real dynamic details, we also add an improved LIF model as the basic neuron dynamic model and record the activation patterns under stimuli. Finally, to prove the network structure matters in brain network communication, we calculate the average physical and network distance between different areas.
2.1 Experimental Settings
The basic methodology of the experiment is described as follows. In this study, a continuous stimulus signal (binary 0-1 signal) is applied to a subset of input neurons (such as visual or olfactory neurons) in the brain. Subsequently, we compute the signal’s propagation process within the brain and record the responses of both intermediate and output neurons. The evaluation criterion is determined by calculating the ratio of neurons that ought to be activated to those that are actually activated in a specific area (see as Equ.26 in appendix B). Given that the brain functions as a type of shallow neural network[53], we document the outcomes at an iteration step size of 5. Moreover, as the input remains constant and no randomization parameters have been introduced, the outcomes of each experiment are consistent, eliminating the need for repetition. For messages in related neurons, please read appendix C.
In linear threshold model, we set $\sigma_{0}=0.8$. In the simulation of the LIF model, numerous hyperparameters need to be used. All the hyperparameters used are listed in the following table. Biological hyperparameters are set according to real biologic experiment test[54]. Current proportion const is the optimal parameter obtained under equidistant search.
Table 1: List of Hyperparameters
Hyperparameter Description Value
$\tau$ Membrane time constant 15.0 ms
$V_{\text{rest}}$ Resting membrane potential -65 mV
$V_{\text{th}}$ Spike threshold -45 mV
[…]
$dt$ time step 0.55 ms
$\alpha$ Current proportion const 0.03/0.045 (visual/olfactory)
2.2 Visualization and Statistic Properties of Drosophila Connectome
@@
In this section we first present the visualization of Drosophila Connectome (Fig.1(a)). Other visualization cases and methods of visualization are presented in appendix F. Visually, the shape of the brain appears to be physically similar, but whether the network structures are similar remains unknown. So we further investigate the network statistics of two hemispheres. We calculated the degree, clustering coefficient, and eigenvector centrality distribution for every node in each hemisphere (Fig.1(d), (e) and (f)). The Pearson correlation coefficient of degree, clustering coefficient, and eigenvector centrality vector in each hemisphere is 0.9986, 0.9988, and 0.9991, respectively. According to Fig.1(d), (e) and (f), these statistical indicators for both hemispheres are extremely similar, indicating that the brain exhibits strong symmetry in network structure.
@@
From Fig.1(c) we can see that the degree, i.e., the number of connections to the nodes in the network, exhibits a long-tail distribution, implying that while a large number of nodes possess few connections, a small number of nodes contains a large number of connections[37]. It implies the real neural network have some same statistic properties as social, electrical and transportation networks. Also, we plot the correlation curves between degree and cell surface area, volume, and cell length (Fig.1). The results indicate that the degree distribution is logarithmically linearly correlated with these characteristics. This logarithmic relationship reveals a fundamental principle in the organization of neural networks in the brain.
To elaborate, as the surface area, volume, or length of a neuron increases, the number of synapses (represented by the degree) it forms grows logarithmically rather than linearly. This non-linear scaling is crucial for maintaining efficient information processing in the brain while preventing an explosion in connectivity that would be computationally and energetically costly.
These characteristics of the brain can explain Weber-Fechner’s law[47]: for any sensory pattern, the perceived intensity is a logarithmic function of the physical intensity. Specifically, the size, length, and area changes can only cause a logarithmic change in the number of synapses the neuron has. This means that as a stimulus becomes stronger, the increase in neural activity or connectivity is proportionally smaller, mirroring the logarithmic perception of intensity described by Weber-Fechner’s law.
(a) Visualization of Drosophila Connectome
(b) Correlation between degree distribution and neuron characteristics
Figure 1: Visualization and statistic properties of Drosophila connectome data set
These statistic results are a valuable supplement to previous findings[48] and are crucial for for improved understanding of the Drosophila brain connectcome structure.
@@ 2.3 Activation Patterns under Different Type of Stimuli
Primarily, to validate the efficacy of our model in signal transmission between input and output neurons, we showcase the activation patterns of intermediate and output neurons throughout the information transmission process. This demonstration serves to substantiate that our model accurately reflects real-world scenarios, shaping a quasi-real activation pattern.
The specific idea for the experiment is following: we stimulate different types of neurons (such as visual input neurons, olfactory neurons, etc.), and record the rate that neurons that are activated in different brain areas. If a specific stimulus causes a significantly greater response than the average in the brain area that it is theoretically supposed to activate, then it is considered that the model can generate a quasi-real activation pattern in the brain. We give 2 types of input neurons (visual, olfactory) stimulus and record the activation patterns of 6 type of neurons: optic (visual), visual_projection (visual), ALLN (olfactory), ALIN (olfactory), ALPN (olfactory) and ALON (olfactory). When we give neuron a stimulus, we set the state of neuron as 1.
To demonstrate the importance of network structure, we add a comparative model that randomly reconnects networks. Specifically, we perform degree-preserving rewiring on a certain proportion of edges in the brain, which means perturbing the network structure while ensuring that the degree of each node remains unchanged. In Fig.2 we set reconnect rate $p=10\%$. For its sensitivity analysis, please read appendix S4.
Fig.2 shows the activation patterns of different models. From Fig. 2(a), (e), and (i), it shows visual stimulus cause a strong impact (significantly more than average) on visual neurons (optic, visual projection) while causing a slight impact (significantly less than average) on other type of neurons under all three kinds of models. Specifically, under visual stimulus, in response to visual stimuli, the optic (vision related area) exhibit a 36.39%, 4.69% and 9.69% ’s activate rates, respectively. At the mean time, no olfactory-related neurons are activated.
In addition, olfactory stimulus cause a strong impact (significantly bigger than average level) on olfactory neurons (ALLN, ALIN, ALPN, ALON), reaching the activation rate by 94.07%, 78.01%, and 69.75%, respectively. And it cause a slight impact (significantly less than average) on other type of neurons. For example, it only activates 0.04% of optic and 0.79% of visual_projection in threshold model (Fig.2(c)).
Although the activation rates of the three models in specific areas differ (due to different activation mechanisms), the significant difference in the activation ratio between correlated and uncorrelated regions indicates the emergence of similar quasi-real activation patterns across all models. This suggests that different activation mechanisms are not the fundamental reason for changes in network functionality.
@@
As for the disturbing models, in Fig.2 (b), (d), (f), (h), (j) and (l), the quasi-real activation patterns totally disappear. Across all activation modes and all types of stimuli, no significant differences in activation regions were observed. For example, in the threshold model, the maximum difference in the activation ratios of all areas under visual stimuli was only 3.162%. Moreover, the range of the regional activation ratios versus the average activation ratio was only 6.32%. Additionally, in the sigmoid and LIF models, no significant activation was observed in any area, indicating that stimuli under network disturbance did not lead to any significant response in any area.
These results indicate that no matter neuron model is, it will lead to similar activation patterns which correspond to reality. After disturbing the network structure, the original quasi-real activation patterns in all models completely disappear. It means the real network structure matters than activation mechanisms in generating the quasi-real activation patterns. Different activation mechanisms are not the primary cause of this realistic activation mechanism; it is the brain’s unique network structure that plays the crucial role. This aligns with the studies of complex systems: individual capabilities are not the determining factors in complex systems; what matters is the organization and interaction structure among individuals. This perspective highlights the importance of individual relationships and the overall system structure in determining emergent phenomena.
(a) threhold, visual
(b) threhold, visual (after disturb)
(c) threhold, olfactory
(d) threhold, olfactory (after disturb)
(e) sigmoid, visual
(f) sigmoid, visual (after disturb)
(g) sigmoid, olfactory
(h) sigmoid, olfactory (after disturb)
(i) LIF, visual
(j) LIF, visual (after disturb)
(k) LIF, olfactory
(l) LIF, olfactory (after disturb)
@@
Figure 2: Experiment of intrinsic neurons’ activation. For disturbing model, we conduct 5 repeated experiments and take the average and standard deviation.

### v4

# Network Structure Governs Drosophila Brain Functionality

@@ Abstract
How intelligence emerges from living beings has been a fundamental question in neuroscience. However, it remains largely unanswered due to the complex neuronal dynamics and intricate connections between neurons in real neural systems. To address this challenge, we leveraged the largest available adult Drosophila connectome data set, and constructed a comprehensive computational framework based on simplified neuronal activation mechanisms to simulate the observed activation behavior within the connectome. The results revealed that even with rudimentary neuronal activation mechanisms, models grounded in real neural network structures can generate activation patterns strikingly similar to those observed in the actual brain. A significant discovery was the consistency of activation patterns across various neuronal dynamic models. This consistency, achieved with the same network structure, underscores the pivotal role of network topology in neural information processing. These results challenge the prevailing view that solely relies on neuron count or complex individual neuron dynamics. Further analysis demonstrated a near-complete separation of the visual and olfactory systems at the network level. Moreover, we found that the network distance, rather than spatial distance, is the primary determinant of activation patterns. Additionally, our experiments revealed that a reconnect rate of at least 1‰ was sufficient to disrupt the previously observed activation patterns. We also observed synergistic effects between the brain hemispheres: Even with unilateral input stimuli, visual-related neurons in both hemispheres were activated, highlighting the importance of interhemispheric communication. These findings emphasize the crucial role of network structure in neural activation and offer novel insights into the fundamental principles governing brain functionality.
@@ 1 Introduction
The neural systems of intelligent beings maintain continuously dynamical communication[1]. This process gives rise to macroscopic emergence across the entire neural system. Such communication fosters the emergence of intelligence in all intelligent life forms. Consequently, comprehending and replicating this process has become an indispensable avenue for unraveling biological intelligence and realizing artificial intelligence[2].
Neural systems are often modeled as interconnected networks[3], with neurons engaging in constant communication via synapses within the network’s topology. To elucidate the relationship between the network itself and its dynamics in a neural network, various models have been developed, spanning from the cellular level to the whole brain. For example, neuronal dynamic models like Leaky Integrate-and-Fire (LIF) model[4] are designed to describe the electrical activity and dynamic behavior of neurons[5]. These models are characterized by the following features: they typically use a set of differential equations to describe the electrophysiological behavior of neurons and usually contain a set of parameters that need to be measured experimentally or estimated. On a larger scale, a vast array of isolated neuronal dynamics, ranging from several to millions, are interconnected and parameter-optimized to simulate the local or global dynamics of the brain[6]. Due to the requirements of commonly used algorithms like backpropagation on network structures, these methods typically do not fully adhere to the structures of real neural networks. Another category of network-based methods applies propagation models on complex networks to neural networks, thereby describing information propagation and emergent mechanisms within neural networks. These methods are collectively referred to as brain network communication models. Such approaches typically involve substantial simplification of neuron models, resulting in fewer parameters, but also leading to the loss of some electrophysiological details of neurons[7]. These two seemingly opposing approaches illustrate that it remains unresolved whether the activation mechanisms of neurons or the connectivity between neurons play the more dominant role. This question is fundamental to understanding how intelligence emerges from neural systems.
Additionally, in recent years, as an innovative way to modify intelligence, artificial neural networks (ANNs) have achieved significant advancements in numerous fields. [8]. Despite these achievements, the structure of ANNs does not strictly mimic the actual neural network structure of biological brains. This discrepancy raises questions about the hypothesis that the inherent structure of a network is crucial for intelligence. Furthermore, the scaling law in the field of machine learning posits that the more parameters an ANN has, the greater the degree of emergent intelligence it can exhibit[9]. These theory hints that the quantity of neurons (parameters) might be a more significant factor in the development of intelligence than the precise arrangement of these neurons in a network, suggesting that the specific network structures found in biological systems may not be essential for the emergence of intelligence.
The observation of neural network structures is crucial for interpreting intelligence from a network perspective. Current neuroimaging techniques, such as Magnetic Resonance Imaging (MRI) and functional MRI (fMRI), offer non-invasive exploration but lack the spatial resolution necessary to resolve individual neurons and synapses. However, these measurement methods face challenges in obtaining information about the positions and connectivity of neurons in the nervous system, making it difficult to understand the nervous system at the neuron-synapse level[10]. Employing electron microscopy (EM) and deep learning techniques for automated neuron reconstruction, brain atlases of a range of animals, including Drosophila larvae[11], have been mapped. A recent study[12] published an adult Drosophila connectome data set, comprising nearly 130,000 neurons and 50,000,000 synapses, with details of neuronal network structure at unforeseen resolution. With full proofreading of the central brain and approximately 80% completion of the optic lobesis presented, the data represents the largest whole-brain connectome of an adult Drosophila to date and provides an unprecedented opportunity for comprehensive network analysis and topological characterization of the Drosophila brain.
Building on the outlined challenges and opportunities, this study utilized the aforementioned Drosophila connectome data set[12] to examine the importance of neuronal network’s structure in simulating brain activation patterns. We constructed a large-scale network communication model framework based on simple mechanisms to simulate activation behavior observed in the connectome. Initially, we established a network representation of the connectome. We then developed a large-scale network communication model where neurons activate upon receiving sufficient stimuli and transmit signals to other neurons. This iterative process generated a global activation pattern. Our analysis employed threshold and sigmoid models as baselines, with a LIF model for comparison. We proposed an evaluation criterion based on the ratio of actual to expected neuron activations.
@@
Additionally, we developed a 3D network visualization tool (see Appendix F), using HTML5’s Canvas and Three.js based on WebGL technology. This software enables real-time observation of each neuron’s communication process in three dimensional space.
@@ 2 Results
2.1 Structural Properties of Drosophila Connectome
The structure of the Drosophila connectome is presented in Fig. 1(a). Visually, the shape of most of the functional regions of the brain appears to be physically similar. From Fig. 1(b) we can see that the degree, i.e., the number of connections to the nodes in the network, exhibits a long-tail distribution, implying that while a large number of neurons possess few synapses, a small number of neurons contains a large number of synapses. For example, more than 20% of the nodes have a degree greater than 50, while only 0.3% of the nodes have a degree greater than 500. This is consistent with previous research on functional brain networks[13].
@@ 2.1 Structural Properties of Drosophila Connectome
To investigate whether the two hemispheres’ network structures are similar in network statistic properties, we conducted a comprehensive network analysis on the two hemispheres. We calculated the degree, clustering coefficient, and eigenvector centrality distribution for every node in each hemisphere. According to Fig. 1(c)-(e), these statistical indicators for both hemispheres are extremely similar, indicating that the brain exhibits strong symmetry in network structure.
@@
Also, we plotted the correlation curves between degree and cell length, cell surface area, and cell size (Fig. 1(f)). The significance coefficients $R$ are 0.85 ($p=0.00$), 0.86 ($p=0.00$), 0.80 ($p=0.00$), respectively. The results indicate that the degree is logarithmically linearly correlated with these characteristics. This logarithmic relationship reveals a fundamental principle in the organization of neural networks in the brain.
Figure 1: Visualization and statistic properties of Drosophila connectome data set. (a) is a visualization of Drosophila connectome. The different colors represent different functional regions. (b) shows the degree distribution of the network. (c), (d) and (e) plot the distribution of degree, clustering coefficient, and eigenvector centrality in each hemisphere, respectively. (f) represents the correlation between degree distribution and neuron characteristics. We use nodes to represent neurons’ cell bodies and use edges to represent the synapses when calculated statistic properties.
@@ 2.2 Activation Patterns under Different Type of Stimuli
In this section, we investigate the impact of different activation mechanisms on network activation patterns by utilizing various network activation mechanisms and observing the changes in activation rates across different areas of the network. Simultaneously, we examine the effect of global perturbations in network structure on activation patterns by altering the network architecture. The optic and visual_projection area are related to vision. ALLN, ALIN, ALON and ALPN are related to olfaction. Other supplemental messages of related neurons are presented in Appendix B.
Fig. 2 illustrates the activation patterns of the large-scale network communication model (LSNC) using threshold, sigmoid, and LIF activation mechanisms. We also evaluated the model’s performance after introducing perturbations. For a detailed description of these models, please refer to Appendix A.2.
From left subgraphs of Fig. 2 (a) and (b), it can be observed that the network model generates quasi-real activation patterns under specific stimuli: It is evident that visual (olfactory) stimuli result the most significant activation on visual (olfactory) neurons such as optic and visual projection neurons (ALLN, ALIN, ALPN, ALON neurons) while activating other neuron types minimally.
Our findings indicate that different activation mechanisms can yield comparable activation patterns. In response to visual stimuli, all three models displayed notable activation in visual-related areas, whereas other areas exhibited minimal activation (left subgraphs of Fig. 2 (a), (c), and (e)). The optic area, for instance, recorded activation rates of 36.39%, 4.69%, and 9.69%. Under olfactory stimuli, the network also exhibits similar activation patterns in olfactory-related areas, with average activation rates of 94.07%, 78.01%, and 69.75%, respectively (left subgraphs of Fig. 2 (b), (d) and (f)). This suggests that the connectivity between neurons is more critical than the activation mechanisms of individual neurons.
@@
However, in the perturbed models (right subgraphs of Fig. 2 (a)-(f)), the original quasi-real activation patterns disappear. No significant differences in activation regions were observed across all activation modes and types of stimuli. For instance, in the model with threshold activation mechanism, the maximum difference of all areas in the activation ratios under visual stimuli was only 3.162%, and the range of regional activation ratios versus the average activation ratio was only 6.32% (right subgraph of Fig. 2(a)), which indicates that the heterogeneity between different regions completely disappears. Additionally, in the sigmoid and LIF models, no significant activation was observed in any area.
Despite variations in activation rates among the three models, the substantial disparity between the activation ratios of relevant and irrelevant areas suggests the emergence of comparable quasi-real activation patterns across all models.
(a) threshold, visual
(b) threshold, olfactory
(c) sigmoid, visual
(d) sigmoid, olfactory
(e) LIF, visual
(f) LIF, olfactory
@@
Figure 2: Comparative test of intrinsic neurons’ activation. For disturbing model, we conduct 5 repeated experiments and take the average and standard deviation. The dotted lines represent average activate rate in brain. In disturbing model, we set reconnect rate $p=20\%$ and $\sigma$ = 0.8.
2.3 Separation of Visual and Olfactory Related Areas
To further investigate the causes of the above observed dynamic pattern that is independent of neuron models, we calculated the synapse count (number of connections) between various regions of the visual and olfactory systems and visualized the results as a heat map (Fig. 3(a)). The findings reveal a significant segregation between the visual and olfactory areas. Specifically, within the visual and olfactory areas, there are 13,309,161 and 882,170 synapses, respectively. However, only 1,837 synapses exist between these two areas. These interconnecting synapses constitute only approximately 0.0138% and 0.208% of the total synapses within the visual and olfactory areas, respectively.
Fig. 3(b) represents the schematic diagram of the connection relationships in the neuronal visual and olfactory systems. We use nodes to represent functional areas and use edges to represent synapses between areas. The thickness of edges represents the number of connections. This diagram indicates that the number of connections within regions is significantly greater than the number of inter-regional connections.
These indicate that the network structure between these two functional systems is almost entirely separate.
Figure 3: Heatmap and schematic diagram illustrating the connections between areas related to vision and olfaction. (a) presents the heatmap, while (b) shows the schematic diagram of the connection relationships.